An automotive manufacturer wants to know the proportion of new car buyers who prefer foreign cars over domestic. Step 1 of 2: Suppose a sample of 822 new car buyers is drawn. Of those sampled, 205 preferred foreign over domestic cars.

Using the data, estimate the proportion of new car buyers who prefer foreign cars. Enter your answer as a fraction or a decimal number rounded to three decimal places.

Step 2 of 2:

Suppose a sample of 822new car buyers is drawn. Of those sampled, 205 preferred foreign over domestic cars. Using the data, construct the 95% confidence interval for the population proportion of new car buyers who prefer foreign cars over domestic cars. Round your answers to three decimal places.

lower endoint= upper endpoint=
Answer #1

part a)

Sample proportion \hat{p} = 205 / 822 = 0.249

b)

95% confidence interval for p is

\hat{p} - Z\alpha/2 * sqrt [ \hat{p} ( 1 - \hat{p} ) / n ] < p < \hat{p} + Z\alpha/2 * sqrt [ \hat{p} ( 1 - \hat{p} ) / n ]

0.249 - 1.96 * sqrt [ 0.249 ( 1 - 0.249) / 822] < p < 0.249 + 1.96 * sqrt [ 0.249 ( 1 - 0.249) / 822]

0.219 < p< 0.279

Lower bound = 0.219

Upper bound = 0.279